# presep= ]e1,e2,e3

### v1

@@ 4.1 Consistency
    $\lim_{n\to\infty}\frac{\overline{F}^{x}(u_{n})+h_{n}^{c_{x}\eta-2}}{nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}))^{2}}=0$, (2).
@@
Assume (1). If $h_{n}\asymp\overline{F}^{x}(u_{n})^{e}$ for $e>(c_{x}(\eta-1)-1)^{-1}$, for $c_{x}\in((\eta-1)^{-1},1)$, and $k_{n}\asymp n^{\delta}$ for $\delta\in(e/(1+e),1)$, then the above limit conditions are satisfied.
@@ 4.2 Asymptotic normality
Assume conditions 1-4, that $F^{x}$ is a continuous distribution and that there exist $c_{x},\alpha_{x}>0,$ and $\beta_{x}>\alpha_{x}$ such that
$\sup_{t\geq 1}t^{\beta_{x}}\left|\overline{F}^{x}(t)-c_{x}t^{-\alpha_{x}}\right|<\infty.$ (7)
@@
$\lim_{n\to\infty}\sqrt{k_{n}h_{n}^{3}}\log\left({n}/{k_{n}}\right)=\lim_{n\to\infty}\sqrt{h_{n}}k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2}}n^{1-\frac{\beta_{x}}{\alpha_{x}}}=0,$ (8)
@@
Assume conditions 1-4, that $F^{x}$ is a continuous distribution and that there exist $c_{x},\alpha_{x}>0,$ and $\beta_{x}>\alpha_{x}$ such that
$\sup_{t\geq 1}t^{\beta_{x}}\left|\overline{F}^{x}(t)-c_{x}\log(t)t^{-\alpha_{x}}\right|<\infty.$ (9)
@@
Assume conditions 1-4, 6, 1, the time series $\left\{\left(X_{j},Y_{j}\right),j\in\mathbb{Z}\right\}$ is stationary and $\alpha$-mixing with $\alpha(j)=\mathcal{O}(j^{-\eta})$, for $\eta>2$, and that for $A>0$ there exists a finite constant $\widetilde{M}^{x}$ not depending on $j$ such that
$\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}\to\widetilde{M}^{x},$ (11)
@@
$\lim_{n\to\infty}\left(h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n})^{-1}\right)=\lim_{n\to\infty}\left(h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n})^{-1/p}\right)=0,$ (12)
@@
In essence, Lemmas 1 and 2 transfer the asymptotic bias conditions into additional speed conditions on the $h_{n}$ and $k_{n}$ sequences. Standard second-order conditions in extremes are often formulated in terms of limits which imply (7); in particular the broad Hall class (Hall and Welsh, 1985) falls into this framework. Second-order conditions akin to (2) are not often encountered in the literature, and our result is inspired by (Kulik and Soulier, 2020, Eq. (9.6.3)). The result of Lemma 3 provides an alternative way of verifying the anti-clustering conditions. For convenience, we now restate the main speed conditions for asymptotic normality of the conditional Hill estimator to hold.
@@
    $\lim_{n\to\infty}\frac{1}{\ell_{n}}=\lim_{n\to\infty}\frac{\ell_{n}}{r_{n}}=\lim_{n\to\infty}\frac{r_{n}}{n}=\lim_{n\to\infty}\frac{n}{r_{n}}\beta_{\ell_{n}}=0$, (3)
@@
    $\lim_{n\to\infty}\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)}{t}\mathrm{d}t=0$, (7).
@@
Assume (1) and that $\left\{(X_{j},Y_{j}),\ j\in\mathbb{Z}\right\}$ is a stationary, $\beta$-mixing time series, with $\alpha$-mixing coefficients $\alpha(j)=\mathcal{O}(j^{-\eta})$, for $\eta>2$, $\beta$-mixing coefficients $\beta(j)=\mathcal{O}(j^{-\nu})$ for $\nu>0$, and that $F^{x}$ is continuous. Assume conditions 1, 3, 6, (11), and the second-order condition (7). Then there exist sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n}),n\in\mathbb{N}\right\}$ such that the first two conditions for consistency, all the above conditions for normality and anti-clustering conditions 2 and 6 are satisfied.
@@
Note that in both of the above theorems, only the first two conditions for consistency are required, by $\beta$-mixing.
@@ 5.1 Bandwidth selection
Various approaches concerning non-parametric regression (see Daouia et al. (2010); Gannoun et al. (2002) ) have proposed a purely data-driven, cross-validation bandwidth selection method
@@
Alternatively, as proposed by Li and Racine (2008), a leave-one-out doubly-smoothed Nadaraya–Watson estimator of the conditional survival function, where also $\left\{Y_{j}\right\}$ is smoothed with distribution function $G$, using bandwidth $b_{n}$ is given by
@@
where $G(v)=\int_{(-\infty,v]}w(u)\mathrm{d}u$, for symmetric, bounded and compactly supported density $w$. In the numeric studies below the kernels are taken to be Gaussian, which deviates from the compact support assumption; this deviation does not pose any practical issues. The asymptotic properties of the bandwidth selection through such leave-one-out procedure for the empirical conditional survival function have, to our knowledge, not been examined in the case of mixing data and are thus material for future work. Presently, we simply mention that this approach provides a global bandwidth independent of both $k_{n}$ and focuses not only on $x$, but on all covariate values at the same time. Since the bandwidth enters the conditional Hill estimator via the kernel on the covariate process $\left\{X_{j}\right\}$, we propose comparing the above bandwidth, which smooths both $\left\{X_{j}\right\}$ and $\left\{Y_{j}\right\}$ with the (global) optimal Sheather-Jones bandwidth, fitted only on the covariate process (see Sheather and Jones (1991)), and the top $k_{n}$ values of $\left\{X_{j}\right\}$, sorted according to the order statistics of $\left\{Y_{j}\right\}$, respectively. We also implement analogous $k_{n}$-dependent bandwidth selection based on only the top order statistics.
@@ 5.2 Simulation study
    Conditional Pareto distribution: Similar construction as the conditional Fréchet, using instead the Pareto distribution function. For a fixed sample size of $n=500$, we compare the (global) optimal Sheather-Jones bandwidth fitted on the entire covariate process, fitted on the top $k_{n}$ values of $\left\{X_{j}\right\}$, sorted according to the order statistics of $\left\{Y_{j}\right\}$ and the method of (13) using (14) as conditional tail estimate.
@@
    Consider a centred Gaussian process W with $\text{var}(W_{j})=\sigma_{j}^{2}$ and Gaussian covariance kernel, and a covariate process $\left\{X_{j}\right\}$. Let $\left\{P^{z}_{i}\right\}$ be the points of a Poisson point process on $(0,\infty)$ having mean measure $\nu_{\alpha_{z}}$, where $\mathbb{R}\ni z\mapsto\alpha_{z}\in(0,\infty)$ is Lipschitz, and let $\textbf{W}^{(i)},i\geq 1$ be i.i.d. copies of W. Then the process $(Y_{j})$ defined by
    $Y_{j}=\bigvee_{i=1}^{\infty}P^{X_{j}}_{i}e^{W_{j}^{(i)}-\alpha_{X_{j}}\sigma_{j}^{2}/2},$ (15)
    is marginally regularly varying with tail index $\alpha_{x}$ on $(X_{j}=x)$ and its construction is shown in Figure 7. We refer to this construction as a Conditional Subordinated Gaussian Max-Stable (CSGMS) process. In the simulation study we let $\left\{X_{j}\right\}$ be i.i.d. standard uniform. To evaluate the sample size, we consider $n=10^{3},10^{4}$, and length-scale $l=0.5,2$ for all $j$, for the low and high dependence cases respectively.
@@
The conditional Fréchet and Pareto are known to satisfy the second-order condition (7). By the remarks on page 224 in Mikosch and Wintenberger (2024), it holds that $V^{(i)}_{j}:=e^{W_{j}^{(i)}-\alpha_{X_{j}}\sigma_{j}^{2}/2}$ in (15) can be represented by $(f_{j}(U_{i}))_{j=1}^{\infty}$ for suitable non-negative measurable functions $f_{j}$ on $(0,1)$, and $\left\{U_{i}\right\}$ are i.i.d. uniform variables on $(0,1)$. By eq. (5.5.7) therein, the CSGMS process above has conditional Fréchet marginals. In particular it also satisfies (7) and we note that all the three constructions satisfy (1). The mapping $z\mapsto\alpha_{z}=1/(3z(z-1)+1)$ is clearly Lipschitz continuous and hence 1 holds. Assumption 6 and equation (11) are verified for the conditional Fréchet and Pareto in Lemma 9, Appendix C, and anti-clustering conditions for the CSGMS hold due to Lemma 10, Appendix C. In the first two cases, the underlying processes are simulated from $\text{AR}(1)$-processes with auto-regressive coefficients absolutely bounded by $1$, which implies $\beta$-mixing with mixing coefficients of exponential decay (see Ibragimov (1962)). When considering the unconditional version of the CSGMS-process, it is possible to show $\beta$-mixing decay of exponential order (see Kulik and Soulier (2020), Example 13.5.4). In (15) we extend the construction by bundling with the covariate sequence, which itself is i.i.d and in particular of exponential $\beta$-mixing decay. Since $z\mapsto\alpha_{z}$ is Lipschitz, then also the bivariate sequence $\left\{X_{j},Y_{j}\right\}$ is of exponential $\beta$-mixing decay. Hence, since the conditional slowly varying functions in (1) are seen to satisfy condition 3 for both Pareto and Fréchet marginals, we may invoke Theorem 8 to conclude that there exists sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n})\right\}$ such that all conditions required in Theorem 5 are met in the above simulation studies.
@@
For the Fréchet setting a), we observe as expected that increasing the dependence structure in the underlying data, results in an increase in finite-sample bias and MSE. In addition, the bias exhibits unstable behaviour when $k_{n}/n$ approaches zero, for the smaller sample size. This highlights the double sub-sampling used in conditional extreme estimators.
For the Pareto setting b), we focus on comparing the bandwidth selection techniques, as this distribution has no second-order component (no bias term). For the low dependence structure, the optimal bandwidth on the upper $k_{n}$ concomitants of $X$, performs slightly better when $k_{n}/n$ is away from zero. However, considering also the high dependence structure, we observe that no technique is superior to another. Hence we use the technique of lowest computational complexity, which is the optimal bandwidth of $X$, for remaining simulations and real-data analysis.
Figure 2: Top panels: Bias. Bottom panels: MSE. We consider the conditional Fréchet model (low dependence (far left) and high dependence (center left)) $n=10^{3},10^{4}$ in solid and dashed, respectively; number of simulations is $N=500$. We also consider the conditional Pareto model (low dependence (center right) and high dependence (far right)) with $n=500,N=200$, and optimal bandwidth of $\left\{X_{j},Y_{j}\right\}$ using least-squares cross validation (solid), optimal bandwidth of $X$ (dashed), optimal bandwidth of upper $k_{n}$ concomitants of $X$ (longdash).
For the CSGMS setting c), we observe a strong positive bias for all $k_{n}/n$. Increasing the sample size dramatically reduces the bias and MSE for the high degree of dependence, to the point where the high dependence scenario catches up to the performance of the low dependence scenario.

### v2

@@ 4.1 Consistency
    $(\overline{F}^{x}(u_{n}^{x})+h_{n}^{c_{x}\eta-2})/(nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}^{x}))^{2})=(k_{n}+nh_{n}^{c_{x}\eta-2})/(h_{n}^{c_{x}}k_{n}^{2})\to 0$, (3).
@@
Assume (1). If $k_{n}\asymp n^{\delta}$ for $\delta\in((\eta+1)^{-1},1)$, $h_{n}\asymp\overline{F}^{x}(u_{n}^{x})^{e}$ for $e\in(0,\delta/(1-\delta)\wedge(\delta(\eta+1)-1)/(2(1-\delta)))$, there exists a $c_{x}\in(0,\infty)$ such that the above limit conditions are satisfied.
@@ 4.2 Asymptotic normality
Assume conditions 1-4, that $F^{x}$ is a continuous distribution and that there exist $c_{x},\gamma(x)>0,$ and $\zeta_{x}\gamma(x)>1$ such that
$\sup_{t\geq 1}t^{\zeta_{x}}\left|\overline{F}^{x}(t)-c_{x}t^{-1/\gamma(x)}\right|<\infty.$ (4)
@@
$\lim_{n\to\infty}\sqrt{k_{n}h_{n}^{3}}\log\left({n}/{k_{n}}\right)=\lim_{n\to\infty}\sqrt{h_{n}}k_{n}^{\zeta_{x}\gamma(x)-\frac{1}{2}}n^{1-\zeta_{x}\gamma(x)}=0,$ (5)
@@
Assume conditions 1-4, that $F^{x}$ is a continuous distribution and that there exist $c_{x},\gamma(x)>0,$ and $\zeta_{x}\gamma(x)>1$ such that
$\sup_{t\geq 1}t^{\zeta_{x}}\left|\overline{F}^{x}(t)-c_{x}\log(t)t^{-\gamma(x)}\right|<\infty.$ (6)
@@
Assume conditions 1-461, the time series $\left\{\left(X_{j},Y_{j}\right),j\in\mathbb{Z}\right\}$ is stationary and $\alpha$-mixing with $\alpha(j)=\mathcal{O}(j^{-\eta})$, for $\eta>2$, and that for $A>0$ there exists a finite constant $\widetilde{M}^{x}$ not depending on $j$ such that
$\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}\to\widetilde{M}^{x},$ (8)
@@
$\lim_{n\to\infty}\left(h_{n}^{1-c_{x}}+h_{n}^{c_{x}(\eta-1)-1}\overline{F}^{x}(u_{n}^{x})^{-1}\right)=\lim_{n\to\infty}\left(h_{n}^{1-c^{\prime}_{x}}+h_{n}^{-1/p+c^{\prime}_{x}(\eta/p-1)}\overline{F}^{x}(u_{n}^{x})^{-1/p}\right)=0,$ (9)
@@
In essence, Lemmas 1 and 2 transfer the asymptotic bias conditions into additional speed conditions on the $h_{n}$ and $k_{n}$ sequences. Standard second-order conditions in extremes are often formulated in terms of limits which imply (4); in particular the broad Hall class (HallWelsh1985) falls into this framework. Second-order conditions akin to (2) are not often encountered in the literature, and our result is inspired by (Kulik2020, Eq. (9.6.3)). The result of Lemma 3 provides an alternative way of verifying the anti-clustering conditions. For convenience, we now restate the main speed conditions for asymptotic normality of the conditional Hill estimator to hold.
@@
    $\lim_{n\to\infty}1/\ell_{n}=\lim_{n\to\infty}\ell_{n}/r_{n}=\lim_{n\to\infty}r_{n}/n=\lim_{n\to\infty}(n/r_{n})\beta_{\ell_{n}}=0$, (3)
@@
    $\sqrt{k_{n}h_{n}}\int_{1}^{\infty}(\mathbb{E}[\widetilde{V}^{x}_{n}(t)]-t^{-1/\gamma(x)}g(x))/t\ \mathrm{d}t\to 0$, (7).
@@
Assume (1) and that $\left\{(X_{j},Y_{j}),\ j\in\mathbb{Z}\right\}$ is a stationary, $\beta$-mixing time series, with $\alpha$-mixing coefficients $\alpha(j)=\mathcal{O}(j^{-\eta})$, for $\eta>2$, $\beta$-mixing coefficients $\beta(j)=\mathcal{O}(j^{-\nu})$ for $\nu>0$, and that $F^{x}$ is continuous. Assume conditions 136, (8), and the second-order condition (4). Then there exist sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n}),n\in\mathbb{N}\right\}$ such that the first two conditions for consistency, all the above conditions for normality and anti-clustering conditions 2 and 6 are satisfied.
@@
Note that in both of the above theorems, only the first two conditions for consistency are required, by $\beta$-mixing. In the simulation study below, we present three non-trivial examples of processes where the above speed conditions and anti-clustering conditions are satisfied.
Remark.
In the proof of Theorem 8, we let $k_{n}\asymp n^{\delta}$, where
$\delta\in\left(0,\ \frac{3}{2}\frac{\nu}{1+\nu}\wedge\frac{2(\zeta_{x}-1/\gamma(x))}{2\zeta_{x}-1/\gamma(x)}\right).$
The restriction $\delta<3\nu/(2(1+\nu))$ is a consequence of allowing the $\beta$-mixing coefficient to decay polynomially of any order. The more common and stronger assumption of exponential decay i.e. $\beta(j)=\mathcal{O}(\rho^{j})$ for $\rho\in(0,1)$ removes this restriction. On the other hand, the restriction $\delta<2(\zeta_{x}-1/\gamma(x))/(2\zeta_{x}-1/\gamma(x))$ is a direct consequence of the required rates in Lemma 1 that arise from the second order assumption (4).
Also notice that the rate of convergence in Theorem 5 is $k_{n}h_{n}$. In the setting of Theorem 8, this rate reduces to
$k_{n}h_{n}\asymp n^{\delta-(1-\delta)e},$
where $\delta$ is in the above interval and $e$ is bounded from below by $\delta/(3(1-\delta))$. To optimize this rate, $\delta$ should be picked as large as possible and $e$ should be as small as possible. For $e$ close to its lower bound, $k_{n}h_{n}$ behaves like $n^{2\delta/3}$. If we consider only the case of exponential decay of the $\beta$-mixing coefficients, the optimal rate is then
$k_{n}h_{n}=o\Big(n^{\frac{2}{3}\frac{2(\zeta_{x}-1/\gamma(x))}{2\zeta_{x}-1/\gamma(x)}}\Big).$
It resembles the rate achieved in the unconditional case under a corresponding second order assumption, $k_{n}=o(n^{\rho})$ where $\rho=2(\zeta_{x}-1/\gamma(x))/(2\zeta_{x}-1/\gamma(x))$ (see Kulik2020, problem 9.6), however the fraction $2/3$ in the exponent is caused by the bandwidth sequence slowing the speed of convergence.
The preprint GoegebeurGuillou2024 obtains similar asymptotic rates for the conditional Hill estimator. We now compare their main assumptions. Their kernel assumptions $(\mathcal{K})$ are similar to ours; they also cover the multidimensional covariate case. They throughout assume a second order condition $(\mathcal{D})$, which is seen to be a special case of equation (4); they do not cover the second order condition of equation (6). Their assumption $(\mathcal{J})$ implies our anti-clustering condition 2. Their strong condition $(\mathcal{X})$ on the joint behaviour of $(X_{0},X_{i})$ is here replaced by simply $g_{X_{0},X_{i}}$ being bounded for every $i$. Their condition (A) is similar to our beta-mixing condition 3 but it depends on an involved probability sequence $\{\nu_{n}\}$. Their condition (B) implies our anti-clustering condition 6. In their Theorem 3.2 and 3.3, the first bandwidth condition corresponds to condition 2, the second to equation (16), the third and fourth to the first and second part of equation (5).
@@ 5.1 Bandwidth selection
In this section we present methods for bandwidth selection, which is one of the main parameters above subject to tuning. The choice of bandwidth should, according to the theorems above, obey certain speed constraints, and the remark in the former section yields that it should depend on the threshold sequence $\{k_{n}\}$. Below we present four selection approaches, two of which are $k_{n}$-dependent and two of which are not. The proposed techniques are assessed in the subsequent subsection.
  1. 1.
    Sheather-Jones bandwidth fitted only on the covariate process (see Sheather_Jones). It’s a plug-in method that provides a global bandwidth, independent of $k_{n}$.
  2. 2.
    Sheather-Jones bandwidth fitted only on the top $k_{n}$ values of $\left\{X_{j}\right\}$, sorted according to the order statistics of $\left\{Y_{j}\right\}$.
  3. 3.
    Various approaches concerning non-parametric regression (see Daouia2010; Gannoun2002 ) have proposed a purely data-driven, cross-validation bandwidth selection method
@@
    Alternatively, as proposed by Li_Racine, a leave-one-out doubly-smoothed Nadaraya–Watson estimator of the conditional survival function, where also $\left\{Y_{j}\right\}$ is smoothed with distribution function $G$, using bandwidth $b_{n}$ is given by
@@
    where $G(v)=\int_{(-\infty,v]}w(u)\mathrm{d}u$, for symmetric, bounded and compactly supported density $w$. In the numeric studies below the kernels are taken to be Gaussian, which deviates from the compact support assumption; this deviation does not pose any practical issues. The asymptotic properties of the bandwidth selection through such leave-one-out procedure for the empirical conditional survival function have, to our knowledge, not been examined in the case of mixing data and are thus material for future work. Presently, we simply mention that this approach provides a global bandwidth independent of both $k_{n}$ and focuses not only on $x$, but on all covariate values at the same time.
  4. 4.
    Finally consider the data-independent case of picking $h_{n}=\sqrt{\log(k_{n})/n}$, which falls into the framework of conditions 241, and equations (5), (16) for certain $\{k_{n}\}$ and appropriately chosen $\{r_{n}\}$.
@@ 5.2 Simulation study
    Conditional Pareto distribution: Similar construction as the conditional Fréchet, using instead the Pareto distribution function. For a fixed sample size of $n=500$, we compare the (global) optimal Sheather-Jones bandwidth fitted on the entire covariate process, fitted on the top $k_{n}$ values of $\left\{X_{j}\right\}$, sorted according to the order statistics of $\left\{Y_{j}\right\}$, the method of equation (10) using equation (11) as conditional tail estimate and the choice $h_{n}=\sqrt{\log(k_{n})/n}$.
@@
    Consider a centred Gaussian process W with $\text{var}(W_{j})=\sigma_{j}^{2}$ and Gaussian covariance kernel, and a covariate process $\left\{X_{j}\right\}$. Let $\left\{P^{z}_{i}\right\}$ be the points of a Poisson point process on $(0,\infty)$ having mean measure $\nu_{1/\gamma(z)}$, where $\mathbb{R}\ni z\mapsto 1/\gamma(z)\in(0,\infty)$ is Lipschitz, and let $\textbf{W}^{(i)},i\geq 1$ be i.i.d. copies of W. Then the process $(Y_{j})$ defined by
    $Y_{j}=\bigvee_{i=1}^{\infty}P^{X_{j}}_{i}e^{W_{j}^{(i)}-(1/\gamma(X_{j}))\sigma_{j}^{2}/2},$ (12)
    is marginally regularly varying with tail index $1/\gamma(x)$ on $(X_{j}=x)$ and its construction is shown in Figure 7. We refer to this construction as a Conditional Subordinated Gaussian Max-Stable (CSGMS) process. In the simulation study we let $\left\{X_{j}\right\}$ be i.i.d. standard uniform. To evaluate the sample size, we consider $n=10^{3},10^{4}$, and length-scale $l=0.5,2$ for all $j$, for the low and high dependence cases respectively.
@@
The conditional Fréchet and Pareto are known to satisfy the second-order condition of equation (4). By the remarks on page 224 in Mikosch_Wintenberger, it holds that $V^{(i)}_{j}:=e^{W_{j}^{(i)}-(1/\gamma(X_{j}))\sigma_{j}^{2}/2}$ in equation (12) can be represented by $(f_{j}(U_{i}))_{j=1}^{\infty}$ for suitable non-negative measurable functions $f_{j}$ on $(0,1)$, and $\left\{U_{i}\right\}$ are i.i.d. uniform variables on $(0,1)$. By eq. (5.5.7) therein, the CSGMS process above has conditional Fréchet marginals. In particular it also satisfies the condition of equation (4) and we note that all the three constructions satisfy equation (1). The mapping $z\mapsto 1/\gamma(z)=1/(3z(z-1)+1)$ is clearly Lipschitz continuous and hence condition 1 holds. Condition 6 and equation (8) are verified for the conditional Fréchet and Pareto in Lemma 9, Appendix C, and anti-clustering conditions for the CSGMS hold due to Lemma 10, Appendix C. In the first two cases, the underlying processes are simulated from $\text{AR}(1)$-processes with auto-regressive coefficients absolutely bounded by $1$, which implies $\beta$-mixing with mixing coefficients of exponential decay (see Ibragimov1962). When considering the unconditional version of the CSGMS-process, it is possible to show $\beta$-mixing decay of exponential order (see Kulik2020, Example 13.5.4). In equation (12) we extend the construction by bundling with the covariate sequence, which itself is i.i.d and in particular of exponential $\beta$-mixing decay. Since $z\mapsto 1/\gamma(z)$ is Lipschitz, then also the bivariate sequence $\left\{X_{j},Y_{j}\right\}$ is of exponential $\beta$-mixing decay. Hence, since the conditional slowly varying functions in equation (1) are seen to satisfy condition 3 for both Pareto and Fréchet marginals, we may invoke Theorem 8 to conclude that there exists sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n})\right\}$ such that all conditions required in Theorem 5 are met in the above simulation studies.
@@
For the Fréchet setting a), we observe that increasing the dependence structure in the underlying data, results in an similar finite-sample bias and MSE. In addition the bias exhibits unstable behaviour when $k_{n}/n$ approaches zero, for the smaller sample size. This highlights the double sub-sampling used in conditional extreme estimators.
For the Pareto setting b), we focus on comparing the bandwidth selection techniques, as this distribution has no second-order component (no bias term). For the both dependence structure, the optimal bandwidth on the upper $k_{n}$ concomitants of $X$, yields slightly lower bias when $k_{n}/n$ is away from zero. For $k_{n}/n$ lower than approximately $0.1$, the data-independent bandwidth $h_{n}=\sqrt{\log(k_{n})/n}$ results in the lowest mean squared error. We observe that neither of these two techniques are superior to another, and hence use the technique of lowest computational complexity, which is $h_{n}=\sqrt{\log(k_{n})/n}$ for the remaining simulations and real-data analysis.
Figure 2: Top panels: Bias. Bottom panels: MSE. We consider the conditional Fréchet model (low dependence (far left) and high dependence (center left)) $n=10^{3},10^{4}$ in solid and dashed, respectively; number of simulations is $N=500$. We also consider the conditional Pareto model (low dependence (center right) and high dependence (far right)) with $n=500,N=200$, and optimal bandwidth of $\left\{X_{j},Y_{j}\right\}$ using least-squares cross validation (long dash), optimal bandwidth of $X$ (dot dash), optimal bandwidth of upper $k_{n}$ concomitants of $X$ (dash) and $h_{n}=\sqrt{\log(k_{n})/n}$ (solid).
For the CSGMS setting c), we observe a strong positive bias for $k_{n}/n$ away from zero for high dependence and low sample size. Increasing the sample size dramatically reduces the bias and MSE for the high degree of dependence, to the point where the high dependence scenario catches up to the performance of the low dependence scenario.